what are the frequency and wavelength of a photon emitted during a transition from n=5 state to then n=2 state in the hydrogen atom

Grade:11

1 Answers

Vikas TU
7 years ago
1/k = R*1^2*(1/2^2 – 1/5^2)
1/k = R(1/4 – 1/25)
     = 21R/100
K = 100/21R
f = c/k = 100c/21R where k is wavelength and c is speed of light